OpenStudy (anonymous):

I need help A 480-kg car moving at 14.4 m/s hits from behind a 570-kg car moving at 13.3 m/s in the same direction. If the new speed of the heavier car is 14.0 m/s, what is the speed of the lighter car after the collision, assuming that any unbalanced forces on the system are negligibly small?

OpenStudy (anonymous):

I took these steps but did not come up with the right answer vf = m2 * v2/m1 = 570*14/480 = 16.63 m/s

OpenStudy (matt101):

This is a conservation of momentum problem. Say M and V are the mass and speed of the heavier car, and m and v are the mass and speed of the lighter car: \[MV_i+mv_i=MV_f+mv_f\]\[(570 \times 13.3)+(480 \times 14.4)=(570 \times 14)+480v_f\] Solving for v(f), you should get 13.6 m/s.

OpenStudy (anonymous):

Great thank you that is the answer